OpenStudy (anonymous):

, : What is the size of a file if it takes 10 ms to download over 1000 km on a 1 Mbps link? Assume speed of light is 2x108 m/s?

OpenStudy (anonymous):

correction speed of light in this problem = 2*10^8 m/s

OpenStudy (anonymous):

Distance = L = 1,000,000m Bandwidth = C = 1,000,000b/s Total Time = T = .010s ( .010s milliseconds? .000010 microseconds) Medium speed = 200,000,000m/s Packet size = X Packet time(Pt) = Packet size / Bandwitdth (Time it takes for the machine to put the file in the link) Travel time(Tt) = Distance / Medium Speed (Time it takes for the packet to travel through the link) Pt = X/C Tt = L/Medium Speed Pt+Tt = 10ms 10ms = X/C + L/Medium Speed X = (10ms - L/Medium Speed) * C X = (.010s - 1,000,000m/200,000,000m/s) * 1,000,000b/s X = (.010s - .005s) * 1,000,000b/s X = (.005s) * 1,000,000b/s X = 5,000 b (considering ms as milliseconds)